# Incorporating Gradients to Rules: Towards Lightweight, Adaptive Provenance-based Intrusion Detection

### v1

@@ V Implementation
The entire system (including data parsing, tag initializing, tag propagating, alarm generating, and training/testing framework) consists of 5KLoC of Python. We will release the code of our system to the public to facilitate further reuse, research, and improvements [^2] https://github.com/LexusWang/CAPTAIN.
We implement the differentiable tag-propagation framework by creating two dictionaries for each tag to store the gradients with respect to $A$ and $G$. As for $T$, because its gradients are unrelated to the node tags, there is no need to store them within the node. The space used to store those gradient dictionaries during training is analyzed in § VI-C2 and § C-B
@@ VI-A1 Datasets
We evaluate Captain using public forensic datasets from the DARPA Transparent Computing program and datasets generated within simulated environments in collaboration with a professional SOC. Table VIII provides an overview of all datasets.
DARPA Datasets. The DARPA Transparent Computing program was organized between 2016 and 2019 to perform several red team assessments. In two weeks, the data collecting teams deploy data collectors on several target hosts [33], which are the target of red teams. We use the public-available datasets from Engagement 3 (E3) and Engagement 5 (E5) in our evaluation. §A-A provides a detailed description of attacks performed in relevant DARPA datasets.
Simulated Environments. Moreover, we collaborate with an industry SOC to acquire additional datasets within realistically simulated scenarios to avoid the problem of “close-world data” [25]. Specifically, the security company furnishes detailed host setups from real-world operating environments, including system configurations and enumerations of active applications. Subsequently, we simulate APT attacks [10] employing the Atomic Red Team framework [6] and leveraging malware sourced from online repositories [23]. The simulated scenarios encompass five APT attacks across three distinct real-world operational settings, elaborated in detail in §A-B.
Data Labeling. For each attack scenario, we label entities and events on the kill chains as malicious and others as benign, according to the provided attack report. Although previous work [45, 34] provided entity-level data labels in their open-sourced repository, they did not specify the data labeling strategy or context. We also observed that the amount of malicious labels in their dataset is excessively large (e.g., over 12 thousand system entities were marked as malicious within a 30-hour period on the Cadets from Engagement 3), which is impractical for a real-world SOC. Thus, we decided to use our data labels in the following evaluation for a fair and unbiased comparison between Captain and the baseline systems.
Note the attack report is written in natural human language, making it challenging to map it to the records in raw auditing logs. Therefore, we make our data labels publicly available to facilitate future research.
@@ VI-A2 Experiment Setup
We deployed Captain and performed all experiments on an Ubuntu 22.04.3 Linux Server with an Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and 1.0 TB memory.
We assume attacks come either from external networks or existing files. Thus, we set their initial integrity score in $A_{0}$ to 0, which means low integrity, in compliance with our conservative settings mentioned in §IV-B.
We partition each dataset into training and testing sets, adhering to the assumption stated in §II-C that the training set should not contain any malicious activities. Specifically, we find out the first attack starting time from the ground truth. Then, all data produced prior to the first attack date becomes the training set, whereas the remaining data becomes the testing set. The duration and size of these training and testing sets are detailed in Table IX in the appendix.
@@ VI-A3 Baseline Detectors
We compare Captain with five SOTA PIDS: Flash [34], Kairos [9], ShadeWatcher [51], NodLink [25], and Morse [19] to evaluate their performance from different perspectives. We chose Flash, Kairos, and NodLink because they are the most recent and SOTA embedding-based PIDS. In addition, their implementations are open-sourced, allowing us to test them on different datasets. Since some other embedding-based PIDS [45, 16, 30] have already been evaluated in [34, 9, 25] and the result shows that Flash, Kairos, and NodLink outperform them, we didn’t include them in our evaluation. We chose Morse since it is the SOTA rule-based PIDS, and we want to evaluate the improvements from our differentiable adaptation framework in Captain. According to [9, 34] and our communication with the authors, the detection systems of ShadeWatcher and ProGrapher are not fully open-source due to proprietary license restrictions. Although the data preprocessing module of ShadeWatcher is open-sourced, we tried our best but could not locate the preprocessing output files used for the following training and testing. We then realized that it was not feasible for us to perfectly replicate their systems for an unbiased comparison. Therefore, we used the code to evaluate the efficiency and latency of data preprocessing and the detection results reported in their paper to evaluate detection accuracy.
For Flash, Kairos, and NodLink, we used their open-sourced code as the basis for the evaluation. We reimplemented Morse according to their paper. All codes used for evaluating the baselines will also be publicly available for future research.
@@ VI-B Detection Accuracy
In this section, we compare Captain with the SOTA PIDS in terms of the detection performance.
@@ VI-B1 False Alarm Events Reduction
We first focus on the reduction of false alarm events, which requires the detector to have the detection granularity at the event level. We compare Captain with the SOTA event-level PIDS ShadeWatcher and the classical tag-based PIDS Morse. Due to the close-source nature of ShadeWatcher, we used the reported detection results in their paper.
The results show that both Morse and Captain successfully detected all attacks in the testing dataset. But Captain reduces the false alarm rate by over 93% (15.66x) compared to ShadeWatcher and over 95% (20.45x) compared to Morse. Additionally, ShadeWatcher holds 80% of events for training and only 10% for testing, while the testing set of Captain is over seven times larger than the testing set of ShadeWatcher. This shows that Captain doesn’t require as much data as ShadeWatcher for training.
@@
One advantage of the rule-based PIDS like Captain and Morse is that it can offer semantic-rich alarms, while ShadeWatcher only flags deviations from patterns observed during training but doesn’t explain the alarm. We then compared the number of false alarms generated by Captain and Morse into different categories. The results are shown in Table IV.
TABLE IV: Comparison with the baseline on different datasets regarding false alarms. We use T-3 and T-5 to represent the Trace datasets from Engagement 3 and Engagement 5, as well as C-3 and C-5 to represent the Cadets datasets from Engagement 3 and Engagement 5. S-1 to S-3 denote SOC-Cloud, SOC-Streaming, and SOC-Dev datasets, respectively.
@@
Table IV shows that Captain can reduce false alarms by over 90% (11.49x) on average for all datasets compared with the non-adaptive Morse. It outperforms Morse in every alarm category. For those remaining false alarms, especially in C-5, S-2, and S-3, Captain cannot remove them by only learning the benign traces for the following reasons. First, some “false” alarms are not purely benign. They are related to some (potential) attack nodes (e.g., some configuration files in /tmp/atScript/atomic-red-team-Gray_dev1.0/* in S-2) but cannot be mapped to specific attack steps directly, so they are not labeled as “malicious.” Second, some nodes and edges don’t exist or trigger any alarm in the training set. Third, since we try to be as conservative as possible, some parameters are tuned just enough to eliminate false alarms in the training set. However, in the testing set, these conservative settings can still cause false alarms again.
@@
Flash 24 29669 29633 10 2705545
Kairos 4 1343 1279 30 2733871
NodLink1 6 38248 38246 0 -
Captain 24 1137 1130 10 2734077
@@
Flash 8 133405 133364 14 176403042
NodLink1 3 4782 4773 0 -
Captain 10 51 47 12 176536396
@@
Flash 2 129859 129800 18 6998697
Kairos 2 3576 3569 18 7124980
NodLink1 4 8743 8735 0 -
Captain 7 2684 2674 13 7125873
@@ VI-B2 False Alarm Entity Reduction
While Captain operates as an event-level detector, we still compare Captain with the SOTA entity-level PIDS Flash, Kairos, and NodLink(Kairos and NodLink give the result at graph-level, but they also support entity-level detection). We transformed the alarm events triggered by Captain into alarm entities using the following method: an entity is considered an alarm entity if it is involved in any alarm events.
@@
It is noteworthy that only a subset of each DARPA dataset is utilized for training and testing in [45, 34, 9]. To simulate the real scenario in SOC, for each DARPA dataset, we utilize the entire dataset for evaluation, employing the method for dividing the training and testing sets as described in §VI-A2. We retrained the model and performed detection using their code. Additionally, we noticed Flash and ThreaTrace did not count the FPs within the two-hop distance from the labeled attack entities in the ground truth while counting the TPs within the two-hop distance from the detected entities [9]. To ensure a fair comparison without providing excessive leniency, we report the TP and FP results, as well as the one-hop FP result, in which we exclude the FP entities within a one-hop distance from the ground truth. The aforementioned experiment settings and data labeling can explain the difference between our results and the results in their papers.
@@
We evaluated the detection accuracy on Cadets, Trace, and Theia datasets of Engagement 3 since those datasets were commonly used by baselines. Table V illustrates the comparison among Captain and the baselines. On all datasets, Captain demonstrates superior performance in identifying more TPs while maintaining fewer FPs. Flash can report a fair amount of TPs but produce an excessive amount of FPs. Flash leverages GNN to learn $k$-hop neighborhood structures. While this technique shows promising performance when there is a significant anomaly within $k$-hop distance, it could degrade when the training set is limited [34]. We observed that although Kairos also detected many anomalous entities, it can filter out FPs with the Anomalous Time Window Queue [9]. It guaranteed decent detection accuracy at the time-window level; the entity-level accuracy, however, is not satisfactory. NodLink can detect all attack processes with a high number of FPs. Moreover, missing other relevant entities, such as files and network sockets, require additional efforts from security analysts to investigate the reported attacks.
@@
Captain successfully detected all attacks in the dataset. Since Captain is an event-level detector, some entities that are not directly related to the attack events will not be reported. For example, if a malware is downloaded and executed, we report the MalFileCreation alarm and the FileExec alarm, which can give the analysts the malware process, file, and parent process. We do not immediately report the entry network node used by attackers to avoid alarm fatigue, which explains the FNs of Captain. The investigations on the relevant nodes can be conducted after the first response.
Another reason for missing TP is that we did not count the attack entities in the 2-hop neighborhood of the detected entities as TP like [34, 9, 45] did. This strict TP calculation method results in more FNs shown in this paper, but we believe it is necessary to assess the performance of PIDS in real-world scenarios.
@@ VI-C1 Detection Latency
In previous sections, we introduced Captain’s rule-based detection framework, highlighting its simplicity compared to other embedding-based PIDS. In this section, we compare Captain with the SOTA PIDS Flash, Kairos, NodLink, and ShadeWatcher to evaluate their response latency. Our latency analysis encompasses three dimensions: buffer time, preprocessing time, and detection time.
We have already introduced buffer time in II-A. Preprocessing time refers to the duration taken to convert the raw audit logs into a data structure that the detection systems can process. It involves data parsing, data cleaning, noise reduction, preprocessing, feature extraction, and so on.
Detection time refers to the interval between the completion of data processing and the moment the result is produced. It is noteworthy that since the detection granularity is different in different PIDS, the detection time might not reflect the actual latency of each system. For instance, a detector based on the whole graph may take longer to deliver detection results of each graph compared with a detector on the entity/event level. However, this does not necessarily mean the former is slower because a large graph could contain many entities and events. Therefore, we calculate the total time spent on detection across the entire dataset.
We use the Trace and Cadets data from DARPA Engagement 3 as the representative datasets for the latency evaluation since they covered all baselines. Since the training stage can be conducted offline, our evaluation focused on these metrics within the context of the testing set. We minimized the bias in our evaluation. For example, Flash divided the data processing into two stages in their code and used files to store intermediate results. And Kairos employed the PostgreSQL database in their code and some files to store the intermediate results. These are the steps not required in a real-time streaming pipeline. Consequently, we excluded the I/O time to focus solely on measuring the ”pure” preprocessing and detection time.
The result is shown in Table VI. Unlike other PIDS that rely on embedding the graph structure and require a buffer time ranging from several minutes to several hours, Captain processes audit logs in a streaming fashion, eliminating the need for any buffer time. Captain is also faster in preprocessing because the logs do not need much preparation for the machine learning model. Because the tags of Captain are much simpler than the state vectors used by Flash and Kairos, and we avoid using text embedding or GNN to aggregate the semantic and contextual graph information, Captain is over 10x faster than the baselines in detection.
@@
1We did not find a clear number in their paper or codes.

### v2

@@ V Implementation
The entire system (including data parsing, tag initializing, tag propagating, alarm generating, and training/testing framework) consists of 5KLoC of Python. We implement the differentiable detection framework by creating two dictionaries for each tag to store the gradients with respect to $A$ and $G$. As for $T$, because its gradients are unrelated to the node tags, there is no need to store them within the node. The space used to store those gradient dictionaries during training is analyzed in § VI-C2 and § E-A
@@ VI-A1 Datasets
We evaluate Captain using public forensic datasets from the DARPA Transparent Computing program and datasets generated within simulated environments in collaboration with an SOC.
DARPA Datasets. The DARPA Transparent Computing program was organized between 2016 and 2019 to perform several red team assessments. In two weeks, the data collecting teams deploy collectors on several target hosts [27]. We use the public-available datasets from Engagement 3 (E3) and Engagement 5 (E5) in our evaluation. [51, 52] provides a detailed description of attacks performed in relevant DARPA datasets.
Simulated Environments. Moreover, we collaborate with an industry SOC to acquire additional datasets within realistically simulated scenarios to avoid the problem of “close-world data” [18]. Specifically, the SOC furnishes detailed host setups from real-world operating environments. Subsequently, we simulate APT attacks [53] employing the Atomic Red Team [54] and online malware repositories [55]. The simulated scenarios encompass five APT attacks across three distinct real-world operational settings, elaborated in detail in §A.
Data Labeling. For each attack scenario, we label entities and events on the kill chains as malicious according to the attack reports. Although previous work [20, 16] provided entity-level data labels, they did not specify the data labeling strategy or context. We also observed that the amount of malicious labels in their dataset is excessively large (e.g., over 12 thousand system entities were marked as malicious within a 30-hour period on the Cadets from DARPA Engagement 3), which is impractical for a real-world SOC. Thus, we decided to use our data labels in the evaluation for a fair and unbiased comparison between Captain and the baseline systems. We also make our data labels publicly available to facilitate future research.
@@ VI-A2 Experiment Setup
We deployed Captain and performed all experiments on an Ubuntu 22.04.3 Linux Server with an Intel(R) Xeon(R) Platinum 8358 CPU @ 2.60GHz and 1.0 TB memory. We partition each dataset into training and testing sets, adhering to the assumption stated in §II-C that the training set should not contain any malicious activities. Specifically, we find the starting time of the first attack. Then, all data produced before that date becomes the training set, whereas the remaining data becomes the testing set.
To avoid biased results due to overfitting, we performed cross-validation when evaluating the performance of Captain. However, standard k-fold cross-validation is unsuitable for streaming logs because the time series cannot be freely split into k groups, as the subsequent tags depend on the previous ones. Therefore, we utilized time series cross-validation, which preserves the chronological order of data. We set a time window and trained the parameters using the data within this window. Afterward, we moved the time window forward until creating k different training sets. In our experiment, we set k=3 and the length of the time windows to be around 2/3 of the total length of the training set. We also discuss overfitting in § VII-A.
Like many other machine learning systems, Captain also relies on appropriate hyperparameters, especially three regularizer coefficients $\alpha,\gamma,\tau$, to train the accurate and robust model (the effect of these hyperparameters are shown in §VI-D2). We provide two methods to fine-tune the hyperparameters, depending on whether a validation set is available. If a validation set exists, we perform grid searching, a generic hyperparameter tuning approach. We train multiple models with different hyperparameter combinations and evaluate their performance on the validation set. The default conservative detector is run on the validation set as the baseline. We select the hyperparameters from the model with the fewest false alarms and equivalent true alarms to the baseline. For instance, the optimal hyperparameters of E3-Cadets is $\alpha=0.1,\gamma=0.1,\tau=0.1$ and the learning rate is 0.01. However, a validation set is usually not available for a PIDS in practice. Even if it exists, it may not encompass all possible attacks. In this case, we propose a heuristic-based method to fine-tune the regularizer coefficients. Recall that the training starts from the conservative adaptive parameters, false alarms in the training set will loosen the conservative settings. The regularizer coefficients are added to the loss function to avoid being too lenient and missing the true alarms. Therefore, the values of the regularizer coefficients $\alpha$, $\gamma$, and $\tau$ are determined by the following question: how many false alarms caused by a node/edge can we tolerate in the training set to avoid being overly lenient? Before training starts, we set a number for the allowed false alarms ($N$), and we can estimate an approximate value of the regularizer coefficients based on $N$: $\alpha\approx 3N,\gamma\approx 3N,\tau\approx 12N$ (see Appendix §D for the mathematical proof). Please note that these are approximate estimations depending on the extent to which we trust the “benign” training data. A practical way might be: set regularizer coefficients using heuristics; if missing alarms occur compared to the conservative baseline, multiply the coefficients by a factor (e.g., 10) and retrain. Conversely, if no alarms are missed for a long time $T$, reduce the coefficients by a factor and retrain.
@@ VI-A3 Baseline Detectors
We compare Captain with five SOTA PIDS: Flash [16], Kairos [15], ShadeWatcher [13], NodLink [18], and Morse [22] to evaluate their performance from different perspectives. We chose Flash, Kairos, and NodLink because they are the SOTA embedding-based PIDS. In addition, their implementations are open-sourced, allowing us to test them on different datasets. Since some other embedding-based PIDS [20, 19, 23] have already been evaluated in these works and the result shows that Flash, Kairos, and NodLink outperform them, we didn’t include them in our evaluation. We chose Morse since it is the SOTA rule-based PIDS, and we want to evaluate the improvements from our differentiable adaptation framework in Captain. To get a fair and unbiased comparison, we use the settings from Morse in the evaluation. According to [15, 16] and our communication with the authors, the detection systems of ShadeWatcher and ProGrapher are not fully open-source due to proprietary license restrictions. Although the data preprocessing module of ShadeWatcher is open-sourced, we tried our best but could not locate the preprocessing output files used for the following training and testing. We then realized that it was not feasible for us to perfectly replicate their systems for an unbiased comparison. Therefore, we used the code to evaluate the efficiency and latency of data preprocessing and the detection results reported in their paper to evaluate detection accuracy. For Flash, Kairos, and NodLink, we used their open-sourced code as the basis for the evaluation. We reimplemented Morse according to their paper.
@@ VI-B1 False Alarm Events Reduction
We first focus on the reduction of false alarm events, which requires the detection granularity at the event level. We compare Captain with the SOTA event-level PIDS ShadeWatcher and the classical rule-based PIDS Morse. Due to the close-source nature of ShadeWatcher, we used the reported detection results in their paper. The results show that both Morse and Captain successfully detected all attacks in the testing dataset. But Captain reduces the false alarm rate by over 93% (15.66x) compared to ShadeWatcher and over 95% (20.45x) compared to Morse. Additionally, ShadeWatcher holds 80% of events for training and only 10% for testing, while the testing set of Captain is over seven times larger than the testing set of ShadeWatcher. This shows that Captain does not require as much data as ShadeWatcher for training.
@@
One advantage of the rule-based PIDS like Captain and Morse is that it can offer semantic-rich alarms, while ShadeWatcher only flags deviations from patterns observed during training without explaining the alarms. We then compared the number of false alarms generated by Captain and Morse into different categories, shown in Table IV.
TABLE IV: Comparison with the baseline on different datasets regarding false alarms. T-3, T-5, C-3, and C-5 represent the Trace and Cadets datasets from Engagement 3 & 5. S-1 to S-3 denote Cloud, Streaming, and Dev datasets from the SOC.
@@
Table IV shows that Captain can reduce false alarms by over 90% (11.49x) on average for all datasets compared with the non-adaptive Morse. It outperforms Morse in every alarm category. Those remaining false alarms, especially in C-5, S-2, and S-3, cannot be removed for the following reasons. First, some “false” alarms are not purely benign. They are related to some (potential) attack nodes (e.g., configuration files in /tmp/atScript/atomic-red-team-Gray_dev1.0/* in S-2) but cannot be mapped to specific attack steps directly, thus not labeled as “malicious.” Second, some nodes and edges do not exist or trigger any alarm in the training set. Third, since our approach aims to be as conservative as possible, some parameters are tuned just enough to eliminate false alarms in the training set. However, these conservative settings may still cause false alarms when applied to the testing set.
@@
Flash 16 4503 4485 10
Kairos 15 1017 1003 11
NodLink1 3 120 114 2
Morse 16 51 43 10
Captain 16 34 26 10
@@
Flash 5 27202 27178 19
NodLink1 4 170 170 0
Morse 10 243 234 14
Captain 10 12 11 14
@@
Flash 2 53230 53050 13
Kairos 12 3566 3422 3
NodLink1 4 62 58 0
Morse 11 220 213 4
Captain 11 194 187 4
@@ VI-B2 False Alarm Entity Reduction
While Captain operates as an event-level detector, we still compare Captain with the SOTA entity-level PIDS Flash, Kairos, and NodLink(Kairos and NodLink give the result at graph-level, but they also support entity-level detection). We transformed the alarm events triggered by Morse and Captain into alarm entities using the following method: an entity is considered an alarm entity if it is involved in any alarm events.
@@
It is noteworthy that only a subset of each DARPA dataset is utilized for training and testing in [20, 16, 15]. To simulate the real scenarios in SOC, we utilize the entire dataset for evaluation, dividing the training and testing sets as described in §VI-A2. We retrained the model and performed detection using their code. Additionally, we noticed Flash and ThreaTrace did not count the FPs within the two-hop distance from the labeled attack entities in the ground truth while counting the TPs within the two-hop distance from the detected entities [15]. To ensure a fair comparison without providing excessive leniency, we report the TP and FP results, as well as the one-hop FP result, in which we exclude the FP entities within a one-hop distance from the ground truth. We also did not count the attack entities in the 2-hop neighborhood of the detected entities as TP like [16, 15, 20] did. These strict experiment settings and data labeling methods can explain the difference between our results and the results in their papers, but we believe it is fair and necessary to assess the performance of PIDS in real-world scenarios.
@@
We evaluated the detection accuracy on Cadets, Trace, and Theia from Engagement 3 since those datasets were commonly used by the baselines. Table V illustrates the comparison among Captain and the baselines. On all datasets, Captain demonstrates superior performance in identifying more TPs while maintaining fewer FPs. Flash can report a fair amount of TPs but produce an excessive amount of FPs. It leverages GNN to learn $k$-hop neighborhood structures. While this technique shows promising performance when there is a significant anomaly within $k$ hops, it could degrade when the training set is limited [16]. Kairos can filter out FPs with the Anomalous Time Window Queue [15]. Although it guaranteed decent detection accuracy at the time-window level; the entity-level accuracy, however, is not satisfactory. NodLink detects most attack processes. However, missing other relevant entities, such as files and network sockets, requires additional expert efforts to investigate the reported alarms. Morse and Captain achieve a lower false positive rate in node-level detection. This is attributed to the semantic-rich alarms that provide additional information for alarm filtering. For instance, a file corruption alarm involves two entities: the process and the corrupted (benign) files. As the corrupted files should not be reported as malicious entities, they can be easily filtered out. This explains why Morse triggers thousands of false alarms at the event level, but only hundreds of nodes are incorrectly alarmed. For embedding-based PIDS, there is no such semantics for alarm filtering. We include the filtered results in Table V as we believe it highlights an advantage of rule-based PIDS.
@@
Captain successfully detects all attacks in the dataset. As an event-level detector, Captain does not report entities that are not directly related to the attack events. For example, if a malware is downloaded and executed, Captain reports the MalFileCreation alarm and the FileExec alarm, which reveals the malware process, file, and parent process. We do not immediately report the entry network node used by attackers to avoid alarm fatigue, which explains the FNs of Captain. The investigations on the relevant nodes can be conducted after the first response.
@@ VI-C Efficiency
In previous sections, we highlighted the simplicity of Captain’s rule-based detection framework compared to other embedding-based PIDS. In this section, we compare Captain with the SOTA PIDS Flash, Kairos, NodLink, ShadeWatcher [^2] Our evaluation was based solely on the open-sourced portion of ShadeWatcher., and Morse to evaluate their efficiency by running their code on our test environment.
@@ VI-C1 Detection Latency
The detection latency analysis encompasses three dimensions: buffer time, preprocessing time, and detection time. We have already introduced buffer time in II-A. Preprocessing time refers to the duration taken to convert the raw audit logs into a data structure that the detection systems can process. It involves data parsing, data cleaning, noise reduction, preprocessing, feature extraction, and so on. Detection time refers to the interval between the completion of data processing and the moment the result is produced. It is noteworthy that since the detection granularity is different in different PIDS, the detection time might not reflect the actual latency of each system. For instance, a detector based on the whole graph may take longer to deliver detection results of each graph compared with a detector on the entity/event level. However, this does not necessarily mean the former is slower because a large graph could contain many entities and events. Therefore, we calculate the total time spent on detection across the entire dataset.
We evaluated the latency on Trace and Cadets from DARPA Engagement 3 since they are covered in the experiments of all baselines. Our evaluation focuses on the testing stage since training can be conducted offline. We modified their code to obtain unbiased results. For example, Flash divides the data processing into two stages in their code and uses files to store intermediate results. And Kairos employs the PostgreSQL database and stores the intermediate results in files. These are the steps not required in a real-time streaming pipeline. Consequently, we exclude the I/O time to focus solely on measuring the “pure” preprocessing and detection time.
The result is shown in Table VI. Unlike other PIDS that embeds the graphs and thus require a buffer time ranging from several minutes to hours, Captain processes audit logs in a streaming fashion, eliminating the need for any buffer time. Captain is also faster in preprocessing because the logs are not preprocessed for the machine-learning model. Moreover, the tags of Captain are much simpler than the state vectors used by Flash and Kairos. By avoiding the use of text embedding or GNNs to aggregate semantic and contextual graph information, Captain achieves detection speeds over 10 times faster than the baseline methods. Lastly, compared to the rule-based PIDS Morse, Captain only takes a few seconds longer to detect logs spanning 4 to 7 days, once again demonstrating the superiority of rule-based PIDS in terms of latency.
@@
1We did not find a clear number in their codes or paper.